A new frame has its length twice it's height.it's made from 126cm of wood what is the length of this frame.
Answer by fractalier(6550)

Call length and height, L and W. (normally it is width)
So from the facts we have
L = 2W and
Perimeter P = 2L + 2W = 126
Substitute the first into the second and get
2(2W) + 2W = 126
4W + 2W = 126
6W = 126
W = 21 cm
L = 42 cm